# Reinforced In-Context Black-Box Optimization

### v2

@@ B.2 Bahavior Algorithms
The datasets are generated with several representatives of heuristic search, EA, and BO as behavior algorithms. We use the implementation in Open Source Vizier [52] for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES. For Hill Climbing and Regularized Evolution, we provide a simple re-implementation. We use the implementation in BoTorch [8] for GP-EI. The details of these behavior algorithms are summarized as follows.
@@
  • Hill Climbing. At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to generate $x_{next}$, which is then evaluated. If $f(x_{next})\textgreater f(x_{best})$, $x_{best}$ is updated to $x_{next}$.
  • Regularized Evolution [50] is an evolutionary algorithm with tournament selection and age-based replacement. We use a population size of $25$ and a tournament size of $5$. At each iteration, a tournament subset is randomly selected from the current population, and the solution with the maximum value is mutated. The mutation operation uniformly selects one of the parameters and mutates it to a random value within the domain.
  • Eagle Strategy [66] without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies. Each firefly emits light whose intensity corresponds to the objective value. At each iteration, for each firefly, a weight is calculated to chase after a brighter firefly and actively move away from darker ones in its vicinity. The position is updated based on the calculated weight.
  • CMA-ES [26] is a popular evolutionary algorithm. At each iteration, candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
  • GP-EI employs GP as the surrogate model and EI [30] as the acquisition function for BO.
@@ B.3 Data
For training, a set of tasks is sampled from the task distribution and the above behavior algorithms are used to collect data. Specifically, for the BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times, except for GP-EI, which is limited to $100$ function samples due to its high time cost. For HPO-B, we use the “v3” meta-training/test splits, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost. All behavior algorithms employ $500$ random seeds in the execution. For the rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times. For testing, we randomly sample $10$ functions for each problem and run each algorithm multiple times to report the average performance and standard deviation.
@@ Appendix C Re-Implementation of OptFormer
OptFormer [14] is a general optimization framework based on transformer [57] and provides an interface to learn policy or function prediction. When provided with textual representations of historical data, OptFormer can determine the next query point $\bm{x}_{t}$, acting as a policy. Additionally, if the context incorporate a possible query point $\bm{x}_{t}$, the framework is capable of predicting the corresponding $y_{t}$, thus serving as a prediction model. We focus on the aspect of policy learning, due to its greater relevance to our work.
@@ Appendix D Generalization
To examine the generalization of our method to unseen function distributions, we conduct experiments to train the model on $4$ of $5$ chosen synthetic functions and test on the remaining one. The results in Figure 5 have shown that RIBBO has a strong generalization ability to unseen function distributions.
@@
We also train the model across all training datasets from BBOB synthetic functions simultaneously, and evaluate on the same functions used in previous sections, specifically Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge. These functions serve as representatives of each category within the BBOB suite. The results are shown in Figure 6. RIBBO demonstrates relatively inferior performance compared to some behavior algorithms, which might be attributed to that the training dataset consists of excessive trajectories from diverse behavior algorithms across different functions, and complicates the process of algorithm identification solely through RTG tokens. Song et al. [53] had similar observation in Section 5.2 of their paper, indicating the significance of textual cues for $\bm{x}$ and the complete failure of the model when using anonymized inputs. However, RIBBO still surpasses both BC and BC Filter, emphasizing the effectiveness of RTG tokens. RIBBO employs RTG tokens to facilitate the identification of diverse algorithms across various functions and to generate sequences with specified regret. Further improvement could be achieved by integrating textual representations for $\bm{x}$ and metadata describing the problem, which may potentially improve the model’s ability to identify and generalize [54]. This leaves as future work.
Figure 6: Generalization by training across the entire BBOB suite.
Appendix E Ablation Studies
@@
Token Aggregation aims to aggregate the information from $(\bm{x}_{i},y_{i},R_{i})$ tokens and establish associations between them. RIBBO employs the Concatenation method (Concat), i.e., concatenating to aggregate $\bm{x}_{i}$, $y_{i}$ and $R_{i}$ to form a single token. This method is compared with two alternative token aggregation methods, i.e., Addition (Add) and Interleaving (Interleave). The addition method integrates the values of each token into one, while the interleaving method addresses tokens sequentially. The results are shown in Figure 7(a). The concatenation method surpasses both the addition and interleaving methods, with the addition method showing the least efficacy. The concatenation method employs a relatively straightforward approach to aggregate tokens, while the interleaving method adopts a more complex process to learn the interrelations among tokens. The inferior performance of the addition method can be due to the summation operator, which potentially eliminates critical details from the original tokens. These details are essential for generating the subsequent query $\bm{x}_{t}$.
@@
Normalization Method is to balance the scales of the function value $y$ across tasks. We compare the employed random normalization with the dataset normalization and the absence of any normalization. Dataset normalization adjusts the value of $y$ by $(y-y_{\rm min}^{i})/(y_{\rm max}^{i}-y_{\rm min}^{i})$ for each task $i$, where $y_{\rm min}^{i}$ and $y_{\rm max}^{i}$ denote the observed minimum and maximum values of $f_{i}$. No normalization uses the original function values directly. The results in Figure 7(b) indicate that random normalization and dataset normalization perform similarly, whereas the absence of normalization significantly hurts the performance. The choice of normalization method directly impacts the computation of RTG tokens and subsequently affects the generation of the desired regret. The lack of normalization leads to significant variations in the scale of $y$, which complicates the training process. Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating both training and inference. However, random normalization can bring additional benefits, such as invariance across various scales of $y$ as mentioned in Wistuba and Grabocka [64] and Chen et al. [14]. Therefore, we recommend using random normalization in practice.
@@
Model Size has an effect on the in-context learning ability. We assess the effects of different model sizes by comparing the performance of the currently employed model size with both a smaller and a larger model. The smaller model consists of $8$ layers, $4$ attention heads, and $128$-dimensional embedding space, while the larger model has $16$ layers, $12$ attention heads, and $384$-dimensional embedding space. The results are shown in Figure 7(c), indicating that the model size has a minimal impact on overall performance. Specifically, the smaller model, due to its limited capacity, shows a reduced performance, while the larger one, potentially more powerful, requires more training data, which can lead to a slight decrease in performance due to overfitting or inefficiency in learning from a limited dataset. This analysis highlights the trade-offs involved in selecting the appropriate model size for optimal performance.
@@
Subsequence Length $\tau$ controls the context length during both the training and inference phases. The process of subsequence sampling acts as a form of data augmentation, enhancing training efficiency. As shown in Figure 7(d), sampling subsequences rather than using the entire history as context, particularly when $\tau=T=150$, leads to improved performance. The computational complexity for causal transformer training and inference scales quadratically with the context length. Therefore, utilizing a shorter $\tau$ can significantly reduce computational demands. However, it is important to note that a shorter $\tau$ might not capture sufficient historical data, potentially degrading the performance due to the insufficient contextual information. This highlights the trade-offs between computational complexity and performance.
(a) Token aggregation
(b) Normalization method
(c) Model size
(d) Subsequence length $\tau$
@@
Appendix F Visualization of Branin Function
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the claims made, and the experimental results reflect the effectiveness of our method.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See the last paragraph of the paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Appendix A and B. The codes are provided in the supplementary.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The codes are provided in the supplementary and the scripts to reproduce experimental results are provided.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setting is presented in Section 4, and the full details are provided in Appendix A and B. The codes are provided in the supplementary.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We show error bars by the length of vertical bars in the figures.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Appendix A.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: See Appendix B.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@ B.2 Bahavior Algorithms
The datasets are generated with several representatives of heuristic search, EA, and BO as behavior algorithms. We use the implementation in Open Source Vizier Song et al. (2022) for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES. For Hill Climbing and Regularized Evolution, we provide a simple re-implementation. We use the implementation in BoTorch Balandat et al. (2020) for GP-EI. The details of these behavior algorithms are summarized as follows.
@@
  • Hill Climbing. At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to generate $x_{next}$, which is then evaluated. If $f(x_{next})>f(x_{best})$, $x_{best}$ is updated to $x_{next}$.
  • Regularized Evolution Real et al. (2019) is an evolutionary algorithm with tournament selection and age-based replacement. We use a population size of $25$ and a tournament size of $5$. At each iteration, a tournament subset is randomly selected from the current population, and the solution with the maximum value is mutated. The mutation operation uniformly selects one of the parameters and mutates it to a random value within the domain.
  • Eagle Strategy Yang & Deb (2010) without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies. Each firefly emits light whose intensity corresponds to the objective value. At each iteration, for each firefly, a weight is calculated to chase after a brighter firefly and actively move away from darker ones in its vicinity. The position is updated based on the calculated weight.
  • CMA-ES Hansen (2016) is a popular evolutionary algorithm. At each iteration, candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
  • GP-EI employs GP as the surrogate model and EI Jones et al. (1998) as the acquisition function for BO.
RIBBO is trained exclusively from offline datasets derived from various source tasks, which are sampled from the task distribution. Each dataset consists of several optimization histories, generated by running a behavior algorithm on a task. No additional assumptions are imposed to the data collection process or the behavior algorithms, thereby permitting the use of any behavior algorithm.
Nevertheless, we can choose appropriate behavior algorithms to help the training. The model is trained on the datasets collected by the behavior algorithms, and the characteristics of these datasets can influence the model’s efficacy. It is advisable to employ well-performing and diverse behavior algorithms to create these datasets. If the datasets are from suboptimal behavior algorithms predominantly, there might be a decline in performance, even with the incorporation of RTG tokens. If the datasets are predominantly produced by behavior algorithms with homogeneous properties, the model may only be adept at addressing specific problems characterized by those properties. Conversely, if the behavior algorithms are diverse, the model can learn the strengths from various algorithms.
@@ B.3 Data
For BBOB and rover problem, a set of tasks is sampled from the task distribution and the above behavior algorithms are used to collect data. Specifically, for the BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times, except for GP-EI, which is limited to $100$ function samples due to its high time cost. For the rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times. For testing, we randomly sample $10$ functions for each problem and run each algorithm multiple times to report the average performance and standard deviation.
For HPO-B, we use the “v3” meta-training/test splits provided by the authors, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost. All behavior algorithms employ $500$ random seeds to collect the training datasets.
@@ Appendix C Re-Implementation of OptFormer
OptFormer Chen et al. (2022) is a general optimization framework based on transformer Vaswani et al. (2017) and provides an interface to learn policy or function prediction. When provided with textual representations of historical data, OptFormer can determine the next query point $\bm{x}_{t}$, acting as a policy. Additionally, if the context incorporate a possible query point $\bm{x}_{t}$, the framework is capable of predicting the corresponding $y_{t}$, thus serving as a prediction model. We focus on the aspect of policy learning, due to its greater relevance to our work.
@@ Appendix D Generalization
We conduct a series of experiments to examine the generalization of our method in this section. We train the model on $4$ of $5$ chosen synthetic functions and test on the remaining one. The results in Figure 5 have shown that RIBBO has a strong generalization ability to unseen function distributions.
@@
We also train the model across all training datasets from BBOB synthetic functions simultaneously, and normalize all the test functions to aggregate results across functions with different output scaling. The results are shown in Figure 6(a), and RIBBO has the best performance.
(a) Aggregation on the entire BBOB
(b) HRR with different $R_{t}$
Appendix E Further Discussions about RTG Tokens
We conduct more discussion about the setting of RTG tokens in this section. The immediate RTG tokens are designed to represent the performance based on cumulative regret over the future trajectories (i.e., $R_{i}=\sum_{t^{\prime}=i+1}^{t}(y^{*}-y_{t^{\prime}})$), rather than focusing solely on instantaneous regret $y^{*}-y_{t+1}$. Thus, setting the immediate RTG to 00 inherently accounts for future regret, thereby enabling the algorithm to autonomously trade-off the exploration and exploitation. We conduct experiments to examine the performance of RIBBO with different values of $R_{t}$. The results are shown in Figure 6(b). Setting to a value larger than 00 will decrease the performance and converge to a worse value, validating the effectiveness of setting the immediate RTG as 00.
Sampling the next query point $\bm{x}_{t}$ is influenced by all the model inputs, including the immediate RTG token $\bm{x}_{t}$ and the previous history $(\bm{x}_{i},y_{i},R_{i})^{t-1}_{i=0}$. The immediate RTG token represents the goal that is intended to be achieved, while the historical data contains information about the problem. These elements are integrated to construct the inputs, influencing the resulting sampled point. Even if the RTG token is set to 00, implying a desire to sample the optimum, the short length of the history suggests limited knowledge about the problem, prompting the model to retain explorative behavior. As the history extends, suggesting a sufficient knowledge of the problem, the preference encoded by the RTG token shifts towards more greedy action. We conduct an experiment about the std of the output Gaussian head of $x$ of RIBBO using GriewankRosenbrock function and the results are shown in Table 2. It is clear that the std is very large in the early stage and becomes small later, which demonstrates the ability to trade-off the exploration and exploitation.
Number of evaluations 1 30 60 90 120 150
std of $\bm{x}$ 0.37 0.17 0.14 0.11 0.03 0.01
Table 2: Std of the output Gaussian head of $x$.
Appendix F Ablation Studies
@@
Token Aggregation aims to aggregate the information from $(\bm{x}_{i},y_{i},R_{i})$ tokens and establish associations between them. RIBBO employs the Concatenation method (Concat), i.e., concatenating to aggregate $\bm{x}_{i}$, $y_{i}$ and $R_{i}$ to form a single token. This method is compared with two alternative token aggregation methods, i.e., Addition (Add) and Interleaving (Interleave). The addition method integrates the values of each token into one, while the interleaving method addresses tokens sequentially. The results are shown in Figure 6(c). The concatenation method surpasses both the addition and interleaving methods, with the addition method showing the least efficacy. The concatenation method employs a relatively straightforward approach to aggregate tokens, while the interleaving method adopts a more complex process to learn the interrelations among tokens. The inferior performance of the addition method can be due to the summation operator, which potentially eliminates critical details from the original tokens. These details are essential for generating the subsequent query $\bm{x}_{t}$.
@@
Normalization Method is to balance the scales of the function value $y$ across tasks. We compare the employed random normalization with the dataset normalization and the absence of any normalization. Dataset normalization adjusts the value of $y$ by $(y-y_{\rm min}^{i})/(y_{\rm max}^{i}-y_{\rm min}^{i})$ for each task $i$, where $y_{\rm min}^{i}$ and $y_{\rm max}^{i}$ denote the observed minimum and maximum values of $f_{i}$. No normalization uses the original function values directly. The results in Figure 6(d) indicate that random normalization and dataset normalization perform similarly, whereas the absence of normalization significantly hurts the performance. The choice of normalization method directly impacts the computation of RTG tokens and subsequently affects the generation of the desired regret. The lack of normalization leads to significant variations in the scale of $y$, which complicates the training process. Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating both training and inference. However, random normalization can bring additional benefits, such as invariance across various scales of $y$ as mentioned in Wistuba & Grabocka (2021) and Chen et al. (2022). Therefore, we recommend using random normalization in practice.
@@
Model Size has an effect on the in-context learning ability. We assess the effects of different model sizes by comparing the performance of the currently employed model size with both a smaller and a larger model. The smaller model consists of $8$ layers, $4$ attention heads, and $128$-dimensional embedding space, while the larger model has $16$ layers, $12$ attention heads, and $384$-dimensional embedding space. The results are shown in Figure 6(e), indicating that the model size has a minimal impact on overall performance. Specifically, the smaller model, due to its limited capacity, shows a reduced performance, while the larger one, potentially more powerful, requires more training data, which can lead to a slight decrease in performance due to overfitting or inefficiency in learning from a limited dataset. This analysis highlights the trade-offs involved in selecting the appropriate model size for optimal performance.
@@
Subsequence Length $\tau$ controls the context length during both the training and inference phases. The process of subsequence sampling acts as a form of data augmentation, enhancing training efficiency. As shown in Figure 6(f), sampling subsequences rather than using the entire history as context, particularly when $\tau=T=150$, leads to improved performance. The computational complexity for causal transformer training and inference scales quadratically with the context length. Therefore, utilizing a shorter $\tau$ can significantly reduce computational demands. However, it is important to note that a shorter $\tau$ might not capture sufficient historical data, potentially degrading the performance due to the insufficient contextual information. This highlights the trade-offs between computational complexity and performance.
(c) Token aggregation
(d) Normalization method
(e) Model size
(f) Subsequence length $\tau$
@@
Appendix G Visualization of Branin Function